PAL2RGB(1) PAL2RGB(1) NAME pal2rgb−convertapalettecolorTIFFimagetoafullcolorimage SYNOPSIS pal2rgb[options]input.tifoutput.tif DESCRIPTION Pal2rgbconvertsapalettecolorTIFFimagetoafullcolorimagebyapplyingthecolormapofthe paletteimagetoeachsampletogenerateafullcolorRGBimage. OPTIONS Optionsthataffecttheinterpretationofinputdataare: −CThisoptionoverridesthedefaultbehaviourofpal2rgbindeterminingwhetherornotcolormapentriescontain16-bitor8-bitvalues.Bydefaultthecolormapisinspectedandif nocolormapentrygreaterthan255isfound,thecolormapisassumedtohaveonly8-bit values;otherwise16-bitvalues(asrequiredbytheTIFFspecification)areassumed.The −Coptioncanbeusedtoexplicitlyspecifythenumberofbitsforcolormapentries:−C 8for8-bitvalues,−C16for16-bitvalues. Optionsthataffecttheoutputfileformatare: −pExplicitlyselecttheplanarconfigurationusedinorganizingdatasamplesintheoutput image:−pcontigforsamplespackedcontiguously,and−pseparateforsamplesstored separately.Bydefaultsamplesarepacked. −cUsethespecificcompressionalgorithmtoencodedimagedataintheoutputfile:−c packbitsforMacintoshPackbits,−clzwforLempel-Ziv&Welch,−czipforDeflate, −cnonefornocompression.Ifnocompression-relatedoptionisspecified,theinputfile’s compressionalgorithmisused. −rExplicitlyspecifythenumberofrowsineachstripoftheoutputfile.Ifthe−roptionis notspecified,anumberisselectedsuchthateachoutputstriphasapproximately8kilobytesofdatainit. BUGS Only8-bitimagesarehandled. SEEALSO tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1